OpenStudy (anonymous):

Find the smallest integer n such that 1 + 2 + 3 + ... + n > 450.

ganeshie8 (ganeshie8):

there is a formula for left hand side : 1 + 2 + 3 +.... + n = n(n+1)/2

ganeshie8 (ganeshie8):

plug that in the given inequality, it becomes to figure out wat n cud be..

ganeshie8 (ganeshie8):

n(n+1)/2 > 450 n(n+1) > 900

ganeshie8 (ganeshie8):

so product two consecutive numbers should give u > 450 do trial and error

ganeshie8 (ganeshie8):

is 29*30 > 900 ? is 30*31 > 900 ? yup corrected :P

OpenStudy (anonymous):

Where did you get the formula?

ganeshie8 (ganeshie8):

http://lmgtfy.com/?q=sum+of+first+n+natural+numbers

OpenStudy (anonymous):

@ganeshie8 I got it. Thank you!

OpenStudy (amistre64):

in general ... let A = 1 + 2 + 3 + ... + n-2 + n-1 + n can be doubled and added together in a special way :) A = 1 + 2 + 3 + ... + n-2 + n-1 + n +A = n + n-1 + n-2 + ... + 3 + 2 + 1 -------------------------------------------- 2A = n+1 + n+1 + n+1 + ... + n+1 + n+1 + n+1 so we have 2A = (n+1), n times. A = n(n+1)/2